Suppose you are undecided to come to work with an electric bike or a car. The distance is 10 (one way) miles. Assume a car will cost 20,000$, the e bike 2500$, suppose you need a loan at 4% interest rate for n 2 years (bike) or 5 years (car) Make and list all the reasonable assumptions on the variable costs and on other possible fixed costs and calculate: 1) the total annual costs (S/y) 2) the cost per mile ($/mile) Are all the relevant factors included? 3) Write some comments on how you would make your decision.
Answer #1

1) Given :- Car Cost = $20000

Electric Bike=$2500

Time = n=2 years (bike) or n=5 years(cars)

Interest rate= 4% p.a

Formula :- Amount *Interest Rate* Time

For Car = 20000*4/100*2= $1600

For Bike=2500*4/100*5=$500

The total annual costs will be

For Car= Interest rate= $1600

For bike = Interest rate=$500

2) For Car The total cost per mile= 20000/1.60934=$12427.45 per mile

For Bike the total cost per mile= 2500/1.60934= $1553.43 per mile

3) According to be electric bike is better because maintenance cost is very less and interest is also economical.